# The Impact of Differential Feature Under-reporting on Algorithmic Fairness

### v1

@@ Abstract.
Predictive risk models in the public sector are commonly developed using administrative data that is more complete for subpopulations that more greatly rely on public services. In the United States, for instance, information on health care utilization is routinely available to government agencies for individuals supported by Medicaid and Medicare, but not for the privately insured. Critiques of public sector algorithms have identified such differential feature under-reporting as a driver of disparities in algorithmic decision-making. Yet this form of data bias remains understudied from a technical viewpoint. While prior work has examined the fairness impacts of additive feature noise and features that are clearly marked as missing, the setting of data missingness absent indicators (i.e. differential feature under-reporting) has been lacking in research attention. In this work, we present an analytically tractable model of differential feature under-reporting which we then use to characterize the impact of this kind of data bias on algorithmic fairness. We demonstrate how standard missing data methods typically fail to mitigate bias in this setting, and propose a new set of methods specifically tailored to differential feature under-reporting. Our results show that, in real world data settings, under-reporting typically leads to increasing disparities. The proposed solution methods show success in mitigating increases in unfairness.
@@ 1. Introduction
Predictive risk models in the public sector have come under criticism over concerns that they are trained on biased data (Mayson, 2019; Richardson et al., 2019; Chouldechova et al., 2018; Berk et al., 2018). In this paper, we consider a specific form of bias. We use the term ‘differential feature under-reporting’ to describe the phenomenon whereby administrative data records are more complete for individuals who have more greatly relied on public services. In the United States, for instance, administrative records often contain medical claims data for those who receive services through public insurance programs (Medicaid / Medicare), but lack information on physical, mental and behavioral healthcare utilization for the privately insured. A lack of recorded medical claims for an individual in this context is often indistinguishable from instances in which no medical claims have been made. Eubanks (2018) writes in her critique of the Allegheny Family Screening Tool (AFST) used in screening child maltreatment referrals,“by relying on data that is only collected on families using public resources, the AFST unfairly targets low-income families for child welfare scrutiny.” We provide a technical analysis of this problem. First, we introduce a statistical model of data collection with differential feature under-reporting. We then present theoretical results that characterize the impact of under-reporting on disparities in selection rates across groups. We describe why standard missing data methods generally fail to mitigate unfairness and, instead, propose new methods based on augmented loss estimation and optimal prediction imputation that are tailored to the under-reporting setting. Lastly, we present empirical results on semi-synthetic and real world data. Our experiments show that, in practice, under-reporting usually leads to increasing disparities and our proposed mitigation methods alleviate this increase.
@@ 2. Background and related work
Missing data literature distinguishes three types of mechanisms: (1) Missing Completely At Random (MCAR) where missing values are independent of both observed and unobserved data, (2) Missing At Random (MAR) where missingness depends on observed variables, and (3) Missing Not At Random (MNAR) where missing values are influenced by systematic factors that are not recorded in the data (Rubin, 1976). In administrative data, records are generally more available for individuals who rely more greatly on public services which often correlates with demographic attributes like race or income levels. Since demographic variables are typically excluded from modeling, this implies an MNAR setting.
@@
In epidemiological surveys, social stigma can lead participants to provide false negative responses (e.g. true $Z$ vs. reported maternal smoking $X$) (King et al., 2001; McKnight et al., 2007; Greenland, 2014; Sechidis et al., 2017). Researchers have proposed various methods to estimate association between $Z$ and outcome $Y$ by leveraging $X$ including correction factors for independence tests (Sechidis et al., 2017; Bross, 1954), adjusted mutual information (Sechidis et al., 2017), odds-ratios (Chu et al., 2006; Edwards et al., 2013; Dosemeci et al., 1990), risk-ratios (Rahardja and Young, 2021; Brenner and Loomis, 1994), and full likelihood approaches (Adams et al., 2019). These methods are typically limited to binary features and outcomes, and focus on statistical inference rather than prediction. Even in settings in which a correctly specified model $f(z)=\mathds{E}\left[{Y\mid Z=z}\right]$ can be estimated using only $X$ and $Y$, it is unclear how to use this model when only under-reported features are available at prediction time. In the study of geographical disease counts, ‘under-reporting’ typically refers to cases with $X\leq Z$. Most work in this area employs Bayesian modeling (Bailey et al., 2005; de Oliveira et al., 2017; Stoner et al., 2019; Gelman et al., 2013) which requires a host of parametric and distributional assumptions as well as informative priors. In single-cell RNA sequencing, ‘zero-inflation’ refers to the phenomenon by which genes are undetected despite being expressed in a cell due to low levels of RNA. Correction methods typically employ Bayesian methods that are highly specialized for the single-cell RNA sequencing task.
@@ 1 $\bm{m}$ 3 1
Features fully observed
Feature values with added random noise $\varepsilon$
Some feature values take default value $m$; $r$ indicates which values are observed
Some feature values take default value $m$; No indicators for missingness
Previous fairness work
No feature mismeasurement
Khani and Liang (2020), Phelps (1972), Aigner and Cain (1977), Chen et al. (2018)
Zhang and Long (2021), Wang and Singh (2021), Jeanselme et al. (2022), Fernando et al. (2021), Fricke (2020), Ahmad et al. (2019)
This work, Eubanks (2018)
@@
Algorithmic fairness literature has studied various types of feature mismeasurement as summarized in Table 1. A commonly studied setting is additive feature noise where, instead of a feature $z_{1}$, we observe a noisy version $x_{1}=z_{1}+\varepsilon$. The random noise $\varepsilon$ is often assumed to be zero-mean, of small variance, and independent of other variables which implies that, while some of the information in the feature is diluted, large portions of the encoded information remains intact. Khani and Liang (2020) show that adding the same amount of feature noise to a group-blind model can introduce statistical loss discrepancy. This is in line with earlier observations from the statistical discrimination literature (Phelps, 1972; Aigner and Cain, 1977). Chen et al. (2018) propose data collection strategies targeted at decreasing discrepancy and come to the conclusion that overcoming differential noise across protected groups may require collection of additional data. In contrast to additive noise, under-reporting removes all information from impacted feature entries and typically biases the feature mean. Some works (e.g. Khani and Liang, 2020) suggest that feature missingness can be modeled as a special case of additive noise by selecting noise terms with very high variance. However, this only covers a special case of under-reporting in which feature entries are missing for all observations.
@@
Feature missingness has been studied in the statistical literature for several decades (e.g. Rubin, 1976). This line of work generally assumes that we observe an indicator variable for missing entries or, equivalently, missing values are clearly marked as such (e.g. as NaN). Various methods with different fairness implications have been proposed to overcome this types of missingness (see also Table 1).
@@
    Imputation. Jeanselme et al. (2022) compare different imputation strategies under clinical presence and find that there is no imputation strategy that reliably outperforms other imputation methods in terms of fairness. Fernando et al. (2021) and Fricke (2020) compare feature imputation to complete case analysis and find that rows with missing values can contribute to fairer outcomes via observed columns. Ahmad et al. (2019) raise concerns about potential harms of imputation in high-stakes settings like recidivism prediction.
@@ 3. Problem setup
Differential under-reporting introduces bias in two ways: (1) Under-reporting in training data influences the estimation of the prediction model (estimation step), and (2) input data with under-reporting leads to biased predictions at test and deployment time (prediction step). We note that, in general, it is not sufficient to recover the true model parameters $\alpha,\beta$ as only biased features are available at prediction time. In fact, our experiments in Section 8 demonstrate that using the true model parameters for prediction with under-reporting can lead to worse fairness outcomes than using a model estimated with biased data directly.
@@
We assume a thresholded prediction setting reminiscent of predictive risk modeling in the public sector. A predictor $f$ is fit on $(X,Y)$ to produce predictions $\hat{Y}=f(X)=f(Z\odot\xi^{G})$. We consider the group-wise shares of predictions that lie above a given threshold $\tilde{y}$: $\mathbb{P}(\hat{Y}\geq\tilde{y}\mid G=g)$ which we refer to as selection rates at threshold $\tilde{y}$. This setup implicitly assumes a predictive risk assessment setting in which the highest risk individuals are selected (e.g. child welfare screenings, fraud detection, federal tax audits). However, it is straightforward to reverse the analysis for scenarios in which low risk leads to selection (e.g. bail decisions in criminal risk assessment). In addition, we assume that being selected is undesirable and the group that is over-selected is disadvantaged. Crucially, this assumption is only made to simplify interpretation of results and we could easily consider the opposite case.
@@
We select the top $C\in[0,1]$ share of the population by thresholding predictions $\hat{Y}$. Given the cumulative distribution function of predictions $F_{\hat{Y}}$, the percentile threshold $C$ implies an absolute threshold $\tilde{y}=F_{\hat{Y}}^{-1}(1-C)$ such that the selection rate for group $g$ can be written as $P(\hat{Y}\geq\tilde{y}\mid G=g)$. In order to isolate the effect of under-reporting, we need to account for a ground truth difference in selection rates. Let $\hat{Y}_{X}$ denote the predictions of a model trained on $(X,Y)$ and $\hat{Y}_{Z}$ the predictions of a model trained on $(Z,Y)$. Since the distributions of $\hat{Y}_{Z}$ and $\hat{Y}_{X}$ generally differ, the predictions imply different thresholds $y^{\prime}=F_{\hat{Y}_{Z}}^{-1}(1-C)$ and $\tilde{y}=F_{\hat{Y}_{X}}^{-1}(1-C)$. With this notation in hand, we are able to define an equity-related metric that allows us to assess the impact of differential feature under-reporting on disparities in selection rates.
@@
In principle, we could directly consider a “difference in difference”: the difference in selection rates between groups $g=0,1$ when selection occurs according to the model $\hat{Y}_{X}$ versus the unbiased predictions $\hat{Y}_{Z}$. However, since we select a fixed share of the population $C$, an increase of the selection rate of group $g$ when moving from $\hat{Y}_{Z}$ to $\hat{Y}_{X}$ already implies a decrease for group $1-g$. It is generally difficult to argue about the excess selection rate $\Delta(g,C)$ analytically. Even in a simple setting with group-dependent Gaussian features $Z\mid G\sim\mathcal{N}\left(\mu_{G},\Sigma_{G}\right)$, there is no closed-form expression for the quantile $y^{\prime}=F_{\hat{Y}_{Z}}^{-1}(1-C)$ and determining the sign of $\Delta(g,C)$ requires analysis of a difference in cdfs which is often intractable. Instead, we simplify the setting and assume that $Z$ follows the same distribution across groups. In this case, the selection rates on the true outcome $Y$ are the same in both groups at every threshold, and we can simplify Definition 1 as follows.
@@
While the majority of our theoretical derivations assume the special case of $Z\perp G$, the empirical portion of this work explores the impact of feature under-reporting on selection rates in the more general setting.
@@ 4. Differential feature under-reporting in linear regression
In this section, we examine the bias that differential feature under-reporting introduces into parameter estimates in linear regression. We consider a setting in which true outcomes are a linear function of the latent feature vectors, i.e. $Y=\alpha+\beta^{T}Z$, which implies that a linear model with access to the true $Z$ recovers the true outcomes $Y$. To simplify notation, we drop the subscripts and write $Y=\hat{Y}_{Z}$ and $\hat{Y}=\hat{Y}_{X}$ for the remainder of the paper. Note that the theoretical portion of this paper considers population-level regressions and we do not discuss finite data results.
Estimates and attentuation bias
@@
where $\Sigma_{XZ}$ denotes the covariance matrix between $X$ and $Z$ and we write $\Sigma_{X}$ for $\Sigma_{XX}$. At first glace, this solution resembles the regression estimates in the more commonly studied additive feature noise case. Assuming $X^{\prime}=Z+U$ where $U$ is independent zero-mean feature noise, we receive $\hat{\beta}=\Sigma_{X^{\prime}}^{-1}\Sigma_{X^{\prime}Z}\beta=(\Sigma_{Z}+\Sigma_{U})^{-1}\Sigma_{Z}\beta$. The factor $\lambda=(\Sigma_{Z}+\Sigma_{U})^{-1}\Sigma_{Z}$ is commonly interpreted as a noise-to-signal ratio and, if $Z$ is one-dimensional, we know that $\mid\hat{\beta}\mid=\lambda\mid\beta\mid<\mid\beta\mid$ which is generally refered to as attentuation bias (e.g. Hausman, 2001; Fuller, 1987). In the under-reporting setting, $\Sigma_{X}$ does not easily separate into terms depending on only the feature or only the mismeasurement. However, in the special case of one-dimensional $Z$ and $Z\perp G$, we can still show that the parameter $\hat{\beta}$ is biased towards zero.
Lemma 1 (Attentuation bias).
Assume the feature $Z$ is one-dimensional and has the same distribution across groups. Then, the least squares regression of $Y$ on the mismeasured feature $X$ yields an estimated slope $\hat{\beta}$ with $\mid\hat{\beta}\mid\leq\mid\beta\mid$.
@@
Real-world prediction settings typically include several features that tend to be correlated in some manner. Assume the feature vector $Z$ is $d$-dimensional, and under-reporting only occurs in the first feature. This means that $X$ coincides with $Z$ in all but the first entry which is computed as $X_{1}=Z_{1}\xi_{1}$ where $\xi_{1}=G\xi_{1}^{1}+(1-G)\xi_{1}^{0}$ and $\xi_{1}^{1}\sim\text{Bern}(m^{0}_{1})$, $\xi_{1}^{0}\sim\text{Bern}(m^{1}_{1})$. We further assume that features $Z_{2},\ldots,Z_{d}$ are uncorrelated and the feature dependence structure is characterized entirely by the correlations between $Z_{1}$ and $Z_{[2:d]}$. In practical settings, this may be achieved via feature orthogonalization. We explicitly exclude cases in which $Z_{1}$ is a perfect linear combination of other features to avoid problems of multicollinearity and assume $\mathds{V}\left[{Z_{k}}\right]>0$ for all $k\in[2:d]$. Given this setting, Proposition A.1 provides a closed form representation for the parameter estimates $\hat{\beta}$. The bias in these estimates can be conceptualized as a generalization of omitted variable bias (Angrist and Pischke, 2008) which is discussed Appendix B. We now assume that $Z\perp G$ which implies that under-reporting is independent of the value that is under-reported. This resembles the assumptions made in previous work on the impact of additive feature noise on fairness (e.g. Khani and Liang, 2020; Phelps, 1972; Aigner and Cain, 1977), and allows us to gain analytical insights that would otherwise remain intractable. First, we examine the behavior of the parameter estimate for the feature with under-reporting, i.e. $\hat{\beta}_{1}$.
@@
    Attentuation bias: $\mid\hat{\beta}_{1}\mid\leq\mid\beta_{1}\mid$.
@@
    Attentuation bias increasing with under-reporting: If under-reporting $1-m_{1}^{g}$ is increasing for one (or both) groups $g\in\{0,1\}$, ceteris paribus, the magnitude of the parameter estimate $\mid\hat{\beta}_{1}\mid$ is decreasing.
@@
[…]
@@
Proposition 2 and 3 tell a compelling story about the effect of under-reporting on parameter estimates in the studied setting. As more feature values default, the regression model places less weight on the mismeasured feature and instead shifts weight to fully observed features with non-zero correlation. This can lead to increasing or decreasing parameter estimates. Analytically, there are cases in which ‘shifting weight’ means that the absolute value of a parameter estimate $\mid\hat{\beta}_{k}\mid$ is decreasing which may appear counterintuitive. For example, consider a setting in which both $Z_{1}$ and $Z_{k}$ have positive true parameters but negative correlation, i.e. $\beta_{1},\beta_{k}>0$ and $\text{Cov}\left[{Z_{1}},{Z_{k}}\right]<0$. In practice, this could occur when there are several mutually exclusive paths to the same outcome. For example, consider prediction of general health risk scores with features including both the number of pediatrician visits in the last year and the number of internist visits. Presumably, these features are negatively correlated because they are relevant for two mutually exclusive parts of the population, i.e. children and adults, but in both columns larger values can be indicative of a high general health risk.

### v2

@@ Abstract.
Predictive risk models in the public sector are commonly developed using administrative data that is more complete for subpopulations that more greatly rely on public services. In the United States, for instance, information on health care utilization is routinely available to government agencies for individuals supported by Medicaid and Medicare, but not for the privately insured. Critiques of public sector algorithms have identified such “differential feature under-reporting” as a driver of disparities in algorithmic decision-making. Yet this form of data bias remains understudied from a technical viewpoint. While prior work has examined the fairness impacts of additive feature noise and features that are clearly marked as missing, little is known about the setting of data missingness absent indicators (i.e. differential feature under-reporting). In this work, we study an analytically tractable model of differential feature under-reporting to characterizethe impact of under-report on algorithmic fairness. We demonstrate how standard missing data methods typically fail to mitigate bias in this setting, and propose a new set of augmented loss and imputation methods. Our results show that, in real world data settings, under-reporting typically exacerbates disparities. The proposed solution methods show some success in mitigating disparities attributable to feature under-reporting.
@@ 1. Introduction
Predictive risk models in the public sector have come under criticism over concerns that they are trained on biased data (Mayson, 2019; Richardson et al., 2019; Chouldechova et al., 2018; Berk et al., 2018). In this paper, we consider a specific form of bias. We use the term ‘differential feature under-reporting’ to describe the phenomenon whereby administrative data records are more complete for individuals who have more greatly relied on public services. In the United States, for instance, administrative records often contain medical claims data for those who receive services through public insurance programs (Medicaid / Medicare), but lack information on physical, mental and behavioral healthcare utilization for the privately insured. A lack of recorded medical claims for an individual in this context is often indistinguishable from instances in which no medical claims have been made. In her critique of the Allegheny Family Screening Tool (AFST) used in screening child maltreatment referrals, Eubanks (2018) writes, “by relying on data that is only collected on families using public resources, the AFST unfairly targets low-income families for child welfare scrutiny.”
We provide a technical analysis of this problem. First, we introduce a statistical model of data collection with differential feature under-reporting. We then present theoretical results that characterize the impact of under-reporting on disparities in selection rates across groups. We describe why standard missing data methods generally fail to mitigate unfairness and, instead, propose new methods based on augmented loss estimation and optimal prediction imputation that are tailored to the under-reporting setting. Lastly, we present empirical results on semi-synthetic and real world data. Our results show that, while in theory under-reporting can decrease disparities, in practice, under-reporting usually leads to increasing disparities and our proposed mitigation methods alleviate this increase.
@@ 2. Background and related work
Missing data literature distinguishes three types of mechanisms: (1) Missing Completely At Random (MCAR) where missing values are independent of both observed and unobserved data, (2) Missing At Random (MAR) where missingness depends on observed variables, and (3) Missing Not At Random (MNAR) where missing values depend on unobserved data (Rubin, 1976). In administrative data, records are more available for individuals who rely more greatly on public services which often correlates with demographic attributes excluded from modeling. This implies an MNAR setting.
@@
In epidemiological surveys, social stigma can lead participants to provide false negative responses (e.g. true maternal smoking status $Z$ vs. reported status $X$) (King et al., 2001; McKnight et al., 2007; Greenland, 2014; Sechidis et al., 2017). Researchers have proposed various methods to estimate association between $Z$ and outcome $Y$ by leveraging $X$ including correction factors for independence tests (Sechidis et al., 2017; Bross, 1954), adjusted mutual information (Sechidis et al., 2017), odds-ratios (Chu et al., 2006; Edwards et al., 2013; Dosemeci et al., 1990), risk-ratios (Rahardja and Young, 2021; Brenner and Loomis, 1994), and full likelihood approaches (Adams et al., 2019). These methods are typically limited to binary features and outcomes, and focus on inferring the relationship between $Y$ and some $Z$ rather than on prediction. In the study of geographical disease counts, (Bailey et al., 2005; de Oliveira et al., 2017; Stoner et al., 2019; Gelman et al., 2013) Bayesian methods have been used to make inference in the under-reporting setting of $X\leq Z$. Such methods require a host of parametric and distributional assumptions as well as informative priors. In single-cell RNA sequencing, under-report arises as ‘zero-inflation,’ which refers to genes going undetected despite being expressed in a cell due to low levels of RNA. Methods correcting for zero-inflation are often Bayesian and highly specialized for the single-cell RNA sequencing task. Note that even if some of the above approaches were applicable in our setting to learn a correctly specified model $f(z)=\mathds{E}\left[{Y\mid Z=z}\right]$ from observations of $X$ and $Y$, it is unclear how to use such models for prediction when only the under-reported features $X$ are available at prediction time.
@@ 1 $\bm{m}$ 3 1
Features fully observed Feature values with added random noise $\varepsilon$ Some feature values take default value $m$; $r$ indicates which values are observed Some feature values take default value $m$; No indicators for missingness
Previous fairness work No feature mismeasurement Khani and Liang (2020), Phelps (1972), Aigner and Cain (1977), Chen et al. (2018) Zhang and Long (2021), Wang and Singh (2021), Jeanselme et al. (2022), Fernando et al. (2021), Fricke (2020), Ahmad et al. (2019) This work, Eubanks (2018)
@@
The algorithmic fairness literature has studied various types of feature mismeasurement as summarized in Table 1. A commonly studied setting is additive feature noise where, instead of a feature $z_{1}$, we observe a noisy version $x_{1}=z_{1}+\varepsilon$. The random noise $\varepsilon$ is often assumed to be zero-mean, of small variance, and independent of other variables. This implies that, while some of the information in the feature is diluted, large portions of the encoded information remains intact. Khani and Liang (2020) show that adding the same amount of feature noise to a group-blind model can introduce statistical loss discrepancy. This is in line with earlier observations from the statistical discrimination literature (Phelps, 1972; Aigner and Cain, 1977). Chen et al. (2018) propose data collection strategies targeted at decreasing discrepancy and come to the conclusion that overcoming differential noise across protected groups may require collection of additional data. In contrast to additive noise, under-reporting removes all information from impacted feature entries and typically biases the feature mean. Some works (e.g. Khani and Liang, 2020) suggest that feature missingness can be modeled as a special case of additive noise by selecting noise terms with very high variance. However, this only covers a special case of under-reporting in which feature entries are missing for all observations.
@@
Feature missingness has been studied in the statistical literature for several decades (e.g. Rubin, 1976). This line of work generally assumes that we observe missingness indicators or, equivalently, missing values are clearly marked (e.g. NaN). Various methods with different fairness implications have been proposed (Table 1).
@@
    Imputation. Jeanselme et al. (2022) compare different imputation strategies under clinical presence and find that there is no imputation strategy that reliably outperforms other imputation methods in terms of fairness. Fernando et al. (2021) and Fricke (2020) compare feature imputation to complete case analysis and find that rows with missing values can contribute to fairer outcomes via observed columns.
@@ 3. Problem setup
Differential under-reporting introduces bias in two ways: (1) Under-reporting in training data influences estimation of the prediction model (estimation step), and (2) input data with under-reporting leads to biased predictions at test time (prediction step). It is generally not sufficient to recover the true model parameters $\alpha,\beta$ as only biased features are available at prediction time. In fact, our experiments in Section 8 demonstrate that using true parameters for prediction can lead to worse fairness outcomes than using a model estimated with biased data.
@@
We assume a thresholded prediction setting reminiscent of predictive risk modeling in the public sector. A predictor $f$ is fit on $(X,Y)$ to produce predictions $\hat{Y}=f(X)=f(Z\odot\xi^{G})$. We consider group-wise shares of predictions above a given threshold $\tilde{y}$: $P(\hat{Y}\geq\tilde{y}\mid G=g)$ which we refer to as selection rates at threshold $\tilde{y}$. This implicitly assumes a setting in which the highest risk individuals are selected (e.g. child welfare screenings, fraud detection, federal tax audits). However, it is straightforward to reverse the analysis for scenarios in which low risk leads to selection (e.g. bail decisions in criminal risk assessment). In addition, we assume that being selected is undesirable. Crucially, this assumption is only made to simplify interpretation and we could easily consider the opposite case.
@@
We assume that the threshold on predictions $\hat{Y}$ is set to achieve a desired overall selection rate ${P}(\hat{Y}\geq\tilde{y})=C\in[0,1]$. Given the cumulative distribution function of predictions $F_{\hat{Y}}$, the percentile threshold $C$ implies an absolute threshold $\tilde{y}=F_{\hat{Y}}^{-1}(1-C)$ such that the selection rate for group $g$ can be written as $P(\hat{Y}\geq\tilde{y}\mid G=g)$. In order to isolate the effect of under-reporting, we need to account for a ground truth difference in selection rates. Let $\hat{Y}_{X}$ denote the predictions of a model trained on $(X,Y)$ and $\hat{Y}_{Z}$ the predictions of a model trained on $(Z,Y)$. When distributions of $\hat{Y}_{Z}$ and $\hat{Y}_{X}$ differ, the predictions imply different thresholds $y^{\prime}=F_{\hat{Y}_{Z}}^{-1}(1-C)$ and $\tilde{y}=F_{\hat{Y}_{X}}^{-1}(1-C)$. With this notation, we define a metric for impact of differential feature under-reporting on disparities in selection rates.
@@
In principle, we could directly consider a “difference in difference”: the difference in selection rates between groups $g=0,1$ when selection occurs according to the model $\hat{Y}_{X}$ versus the unbiased predictions $\hat{Y}_{Z}$. However, since we select a fixed share of the population $C$, an increase of the selection rate of group $g$ when moving from $\hat{Y}_{Z}$ to $\hat{Y}_{X}$ already implies a decrease for group $1-g$. It is generally difficult to argue about the excess selection rate $\Delta(g,C)$ analytically. Even in a simple setting with group-dependent Gaussian features $Z\mid G\sim\mathcal{N}\left(\mu_{G},\Sigma_{G}\right)$, there is no closed-form expression for the quantile $y^{\prime}=F_{\hat{Y}_{Z}}^{-1}(1-C)$ and determining the sign of $\Delta(g,C)$ requires analysis of a difference in cdfs which is often intractable. Instead, we simplify the setting and assume that $Z$ follows the same distribution across groups. In this case, the selection rates on the true outcome $Y$ are the same in both groups at every threshold, and we can simplify.
@@
While the majority of our theoretical derivations assume the special case of $Z\perp G$, the empirical portion of this work explores the impact of feature under-reporting in the more general setting. To clarify which assumptions are sufficient for which finding in the paper, we supply a summary table in Appendix A, alongside descriptions in the main text.
@@ 4. Differential feature under-reporting in linear regression
In this section, we examine the bias that differential feature under-reporting introduces into regression parameter estimates. We consider a setting in which true outcomes are a linear function of latent features, i.e. $Y=\alpha+\beta^{T}Z$, which implies that a linear model with access to the true $Z$ recovers the true outcomes $Y$. Dropping subscripts, we write $Y=\hat{Y}_{Z}$ and $\hat{Y}=\hat{Y}_{X}$. Note that this section focuses on population-level regression.
Estimates and attenuation bias
@@
where $\Sigma_{XZ}$ denotes the covariance matrix between $X$ and $Z$ and we write $\Sigma_{X}$ for $\Sigma_{XX}$. At first glace, this solution resembles the regression estimates in the more commonly studied additive feature noise case. Assuming $X^{\prime}=Z+U$ where $U$ is independent zero-mean feature noise, we obtain $\hat{\beta}=\Sigma_{X^{\prime}}^{-1}\Sigma_{X^{\prime}Z}\beta=(\Sigma_{Z}+\Sigma_{U})^{-1}\Sigma_{Z}\beta$. The factor $\lambda=(\Sigma_{Z}+\Sigma_{U})^{-1}\Sigma_{Z}$ is commonly interpreted as a noise-to-signal ratio and, if $Z$ is one-dimensional, we know that $|\hat{\beta}|=\lambda|\beta|<|\beta|$ which is generally referred to as attenuation bias (e.g. Hausman, 2001; Fuller, 1987). In the under-reporting setting, $\Sigma_{X}$ does not easily separate into terms depending on only the feature or only the mismeasurement. However, in the special case of one-dimensional $Z$ and $Z\perp G$, we can still show that the parameter $\hat{\beta}$ is biased towards zero.
Lemma 1 (Attenuation bias).
Assume the feature $Z$ is one-dimensional and has the same distribution across groups. Then, the least squares regression of $Y$ on the mismeasured feature $X$ yields an estimated slope $\hat{\beta}$ with $|\hat{\beta}|\leq|\beta|$.
@@
Real-world prediction settings typically include multiple, often correlated, features. Assume the feature vector $Z$ is $d$-dimensional, and under-reporting only occurs in the first feature, $Z_{1}$. This means that $X$ coincides with $Z$ in all but the first entry which is computed as $X_{1}=Z_{1}\xi_{1}$ where $\xi_{1}=G\xi_{1}^{1}+(1-G)\xi_{1}^{0}$ and $\xi_{1}^{1}\sim\text{Bern}(m^{0}_{1})$, $\xi_{1}^{0}\sim\text{Bern}(m^{1}_{1})$. We further assume that features $Z_{2},\ldots,Z_{d}$ are uncorrelated and the feature dependence structure is characterized entirely by the correlations between $Z_{1}$ and $Z_{[2:d]}$. This assumption is without loss of generality because we can always apply orthogonalization to the features. We explicitly exclude cases in which $Z_{1}$ is a perfect linear combination of other features to avoid problems of multicollinearity and assume $\mathds{V}\left[{Z_{k}}\right]>0$ for all $k\in[2:d]$. Given this setting, Proposition B.1 provides a closed form representation for the parameter estimates $\hat{\beta}$. The bias in these estimates can be conceptualized as a generalization of omitted variable bias (Angrist and Pischke, 2008) which is discussed Appendix C. We now assume that $Z\perp G$ which implies that under-reporting is independent of the value that is under-reported. This resembles the assumptions made in previous work on the impact of additive feature noise on fairness (e.g. Khani and Liang, 2020; Phelps, 1972; Aigner and Cain, 1977), and allows us to gain analytical insights that would otherwise remain intractable. First, we examine the behavior of the parameter estimate for the feature with under-reporting, i.e. $\hat{\beta}_{1}$.
@@
    Attenuation bias: $|\hat{\beta}_{1}|\leq|\beta_{1}|$.
@@
    Attenuation bias increasing with under-reporting: If under-reporting $1-m_{1}^{g}$ is increasing for one (or both) groups $g\in\{0,1\}$, ceteris paribus, the magnitude of the parameter estimate, $|\hat{\beta}_{1}|$, is decreasing.
@@
    Shifting weight: If under-reporting $1-m_{1}^{g}$ is is increasing for one (or both) groups $g\in\{0,1\}$, ceteris paribus, $\hat{\beta}_{k}$ is increasing if $\text{sign}\left(\beta_{1}\text{Cov}\left[{Z_{1}},{Z_{k}}\right]\right)=+1$, and decreasing if $\text{sign}\left(\beta_{1}\text{Cov}\left[{Z_{1}},{Z_{k}}\right]\right)=-1$.
@@
Proposition 2 and 3 tell a compelling story about the effect of under-reporting on parameter estimates in the studied setting. As more feature values default, the regression model places less weight on the mismeasured feature and instead shifts weight to fully observed features with non-zero correlation. This can lead to increasing or decreasing parameter estimates. Analytically, there are cases in which ‘shifting weight’ means that the magnitude of a parameter estimate, $|\hat{\beta}_{k}|$, is decreasing, which may appear counterintuitive. For example, consider a setting in which both $Z_{1}$ and $Z_{k}$ have positive true parameters but negative correlation, i.e. $\beta_{1},\beta_{k}>0$ and $\text{Cov}\left[{Z_{1}},{Z_{k}}\right]<0$. In practice, this could occur when there are several mutually exclusive paths to the same outcome. For example, consider prediction of general health risk scores with features including both the number of pediatrician visits in the last year and the number of internist visits. Presumably, these features are negatively correlated because they are relevant for two mutually exclusive parts of the population, i.e. children and adults, but in both columns larger values can be indicative of a high general health risk.